OpenStudy (anonymous):

35. What is the probability of rolling a sum of 8 on at least one of two rolls of a pair of number cubes?

OpenStudy (anonymous):

its the number of ways of getting 8 out of all the numbers on 2 dice. There are 36 numbers on two dice and 6 ways of getting an 8 so the probability is 6/36 + 6/36 because we are thowing the pair twice this gives 12/36 = 1/3

OpenStudy (zarkon):

there are 5 ways to get an 8

OpenStudy (anonymous):

six if 4 and 4 are not chosen as a repeat.

OpenStudy (zarkon):

there is only one way to get a 4 and 4...that is you get a 4 then a 4

OpenStudy (anonymous):

Ok well the rest is still good just 5/36 instead

OpenStudy (zarkon):

it is not...you are missing a little bit

OpenStudy (zarkon):

P( A or B)=P(A)+P(B)-P(A and B) or use the complement P(A or B)=1-P(A' and B')

OpenStudy (anonymous):

but hold on just looked it up cos I was confused and your formula is for events that are not mutually exclusive like pulling a jack and a spade from a deck. These events are exclusive so P(A)+P(B) is the correct probability in this case.

OpenStudy (zarkon):

they are not mutually exclusive...they are independent

OpenStudy (zarkon):

the formulas I gave always hold true. you can combine them with independence to get P(A or B)=P(A)+P(B)-P(A and B)=P(A)+P(B)-P(A)P(B) or P(A or B)=1-P(A' and B')=1-P(A')P(B')

OpenStudy (zarkon):

\[P(A)+P(B)-P(A)P(B)=\frac{5}{36}+\frac{5}{36}-\frac{5}{36}\frac{5}{36}=\frac{335}{1296}\] or \[1-P(A')P(B')=1-\frac{31}{36}\frac{31}{36}=\frac{335}{1296}\]

OpenStudy (zarkon):

this also follows a binomial distribution \[\sum_{x=1}^{2}{2\choose x}\left(\frac{5}{36}\right)^x\left(1-\frac{5}{36}\right)^{2-x}=\frac{335}{1296}\]